### v2

# AI-Driven Grading and Moderation for Collaborative Projects in Computer Science Education

Abstract
Collaborative group projects are integral to computer science education, as they foster teamwork, problem-solving skills, and industry-relevant competencies. However, assessing individual contributions in group settings has long been challenging. Traditional assessment strategies, such as the equal distribution of grades or subjective peer assessments, often fall short in terms of fairness, objectivity, and scalability, particularly in large classrooms. This paper introduces a semi-automated, AI-assisted grading system that evaluates both project quality and individual effort using repository mining, communication analytics, and machine learning models. The system comprises modules for project evaluation, contribution analysis, and grade computation, and integrates seamlessly with platforms such as GitHub. A pilot deployment in a senior-level course demonstrated high alignment with instructor assessments, increased student satisfaction, and reduced instructor grading effort. We conclude by discussing implementation considerations, ethical implications, and proposed enhancements to broaden applicability.
I Introduction
Group projects are foundational to computer science curricula, providing students with hands-on experience in software engineering practices, including Agile methodologies, version control, collaborative coding, and iterative testing. These collaborative endeavors mirror professional environments and are invaluable in preparing students for real-world development workflows. However, assigning fair individual grades remains difficult, particularly when group sizes are large or contributions are unevenly distributed.
Instructors often lack granular visibility into team dynamics, leading to assessments based on incomplete information. Peer evaluations, though valuable, can be influenced by bias, social dynamics, and non-academic considerations. Manual oversight becomes unfeasible at scale, prompting the need for automated tools that provide accurate, transparent, and scalable assessments.
The core objectives of this research are to design, implement, and evaluate an AI-based grading and moderation system that (1) evaluates overall group project quality using standardized and customizable metrics, (2) quantifies individual contributions using objective data sources like version control and communication logs, (3) generates fair and explainable individual grades, (4) reduces instructor workload while maintaining oversight and flexibility, and (5) enhances transparency and fairness, thereby improving student satisfaction.
II Related Work
Traditional grading methods include equal grading (where all group members receive the same grade regardless of their individual contributions), peer evaluation tools, and instructor observation. Peer evaluation systems, such as CATME and SPART+, help assess team member performance from both social and task-based perspectives. However, these systems are susceptible to bias, collusion, and peer pressure [1, 9]. Additionally, peer reviews may fail to reflect real technical contributions, especially in large or distributed teams [10].
Instructor observation attempts to gauge individual effort based on limited interactions, project meetings, or code reviews. Yet, this method is time-consuming, not scalable, and lacks consistency, particularly in large classes or remote settings [2, 11].
To overcome these limitations, recent research has explored the use of software repositories to estimate contributions. Metrics such as number of commits, lines changed, file ownership, and authorship of specific code blocks offer quantifiable proxies for student effort [3, 4, 12, 13]. However, committing frequency or volume alone does not always correlate with meaningful contributions, as students may make frequent but trivial changes (e.g., whitespace edits or repeated formatting). Some tools, such as GitGrade, Codequiry, and Gitorium, provide basic metrics from version control logs but lack integrated evaluation frameworks that account for quality, context, or collaboration patterns. More advanced solutions utilize Natural Language Processing (NLP) to extract information from issue trackers, pull request discussions, and documentation comments [5, 14, 15]. This yields a more holistic view of individual contributions, capturing leadership roles, communication, and task complexity—not just code volume.
The broader field of Artificial Intelligence in Education (AIEd) applies machine learning, educational data mining, and analytics to support adaptive instruction, feedback, and engagement monitoring [6]. However, automated grading systems are criticized for a lack of transparency, limited adaptability, and reliance on historical patterns, which may disadvantage students with atypical learning styles.
Finally, ethical considerations, including algorithmic fairness, data privacy, surveillance, and student consent, are critical [8]. These must be carefully addressed before integrating AI into high-stakes assessments, such as course grades and other evaluations.
III System Design and Architecture
In this paper, we propose an AI grading system comprising three core modules: the Project Quality Assessment Module (PQAM), the Individual Contribution Analyzer (ICA), and the Grading Engine (GE). We will illustrate each one and explain how they interact in the following sections.
III-A Project Quality Assessment Module (PQAM)
This module is a comprehensive evaluation component that assesses both the technical and qualitative integrity of a group software project. It leverages a range of automated tools and AI models to ensure the project meets academic and industry standards across multiple dimensions. The module supports objective grading, promotes best practices, and reduces instructors’ manual workload. This module contains five submodules: Code Quality, Testing Coverage, Documentation, Functionality, and Usability. The final output of this module is a general Project Quality Score (PQS) that represents the overall quality of the group’s project. Each module, by default, holds the same weight. We can make weights configurable per project based on learning outcomes and project context.
III-A1 Code Quality
The Code Quality submodule uses static analysis to examine the source code without executing it. The analysis identifies structural flaws, maintainability issues, and stylistic violations. It first builds Complexity Metrics that include (1) Cyclomatic Complexity, which measures the number of independent paths through the code. A high value suggests complex, harder-to-maintain code, and (2) Halstead Metrics, which quantifies computational complexity based on operators and operands. These values provide insight into code readability, maintainability, and error proneness.
Second, it uses Style Adherence to verify code compliance with specific programming-language standards. For example, we can use Python tools such as flake8, pylint, and black to verify compliance with PEP8 standards. For JavaScript, tools such as ESLint and Prettier ensure that code adheres to established best practices and organizational standards.
Third, it uses Code Duplication Detection to apply tools such as SonarQube, PMD, or jscpd, which detect copy-pasted code blocks, indicating poor abstraction and reusability practices.
III-A2 Testing Coverage
The Testing Coverage submodule is crucial for project reliability, as robust testing practices assess the extent to which tests cover the codebase. The available coverage tools for Python include pytest-cov and coverage.py, which can generate online reports, including branch and path coverage. For Java, JaCoCo or JUnit integration helps assess method and class-level coverage. For JavaScript, Istanbul/NYC, or Jest produces detailed test coverage maps. Furthermore, beyond simply counting lines covered, some AI models or advanced scripts can analyze whether edge cases, exception handling, and input validation are adequately tested to enhance coverage depth.
III-A3 Documentation
The Documentation submodule reflects a team’s ability to communicate and maintain their software. This part uses NLP (Natural Language Processing) techniques to assess the quality and relevance of project documentation. For readability and clarity, we can use NLP models such as BERT, spaCy, or GPT-based classifiers to evaluate the grammar, structure, and coherence of README files and manuals. For completeness checks, we can utilize heuristics and AI models to verify the inclusion of essential elements, such as installation steps, usage instructions, architecture overview, contributor guidelines, and licensing. For inline comments and docstrings, we can use parser tools such as Docformatter, Doxygen, or Sphinx to assess whether functions/classes are documented, or perform semantic similarity analysis to determine whether the comment aligns with the code logic or is superficial. For project Wikis or notebooks, such as Markdown or Jupyter Notebooks, Notebook-based documentation is evaluated for its technical depth, clarity of thought, and instructional value.
III-A4 Functionality
The Functionality submodule is a direct evaluation of whether the software works as intended. We can enable an Automated Test Execution process that executes predefined or student-supplied unit and integration tests in sandboxed environments (e.g., Docker containers). The grading system compares actual outputs with expected outcomes and flags deviations. If students use tools like GitHub Actions, Travis CI, or GitLab CI, the system checks for passing build statuses and continuous testing practices. Moreover, feature validation can be performed using AI-based log analysis or semantic parsing of the project requirements to verify that the implemented functionality aligns with the deliverables.
III-A5 Usability
The Usability submodule assesses the system’s intuitiveness, responsiveness, and accessibility, particularly when the project includes a GUI or web-based interface. Several major components can be used to evaluate code usability. For example, heuristic evaluation (optional/manual augmentation) can apply usability heuristics (e.g., Nielsen’s 10 principles) to detect issues like inconsistent navigation, lack of feedback, or poor error messaging. UI testing frameworks such as Selenium, Cypress, or Puppeteer can be used to run scripted tests across various devices and browsers. For accessibility checks, Lighthouse, Axe-Core, or Pa11y can be integrated to verify WCAG (Web Content Accessibility Guidelines) compliance. Color contrast, screen reader compatibility, and keyboard navigation are some of the key elements analyzed. For responsiveness analysis, UI components are tested across a range of screen resolutions to verify compatibility with both mobile and desktop devices.
In general, the Project Quality Assessment Module (PQAM) (as shown in Fig. 1) serves as a holistic evaluation engine that combines static code analysis, test validation, NLP-based documentation review, and UI/UX assessments. It fosters best engineering practices and enables fair, scalable, and transparent grading of collaborative software projects in computer science education.
Code Quality Test Coverage Documentation Functionality Usability Project Quality Score
Figure 1: Structure of the Project Quality Assessment Module (PQAM), highlighting submodules contributing to the Project Quality Score.
III-B Individual Contribution Analyzer (ICA)
This analyzer is a core module that objectively assesses each student’s involvement in a collaborative group software project. Traditional group assessments often obscure individual effort, potentially leading to unfair grading. ICA tackles this by analyzing multi-faceted signals from the development lifecycle to quantify and qualify each student’s contribution. ICA operates across four key dimensions: Commit History Analysis, Code Ownership, Issue Tracker Participation, and Code Review Evaluation.
III-B1 Commit History Analysis
Commit History Analysis evaluates version control activity, focusing on meaningful code contributions rather than superficial edits. First, it filters trivial changes. AI heuristics and regular expression (regex) rules filter out non-substantive changes, such as whitespace adjustments, minor comment updates, or variable renaming. This prevents students from artificially inflating their contribution through “commit spam.” Second, for temporal and behavioral analysis, the system can observe patterns such as commit frequency, time of day, and burst activity (indicating rushed contributions). Clustering algorithms (e.g., DBSCAN) can be used to distinguish between organic and last-minute participation. Third, quantitative metrics, including line additions, deletions, and net code changes, are computed. Furthermore, special weighting can be applied to the commits involving test cases, documentation, or refactors.
III-B2 Code Ownership
Code Ownership attributes authored code lines to individuals by leveraging the git blame command. In more detail, every line of code in the final submission is mapped to its author using Git history. Edited lines are tracked across commit history to identify both original authors and significant editors. Full ownership is assigned to those who write or significantly revise code logic. Partial credit is given for minor edits or formatting adjustments. Code ownership is aggregated at the function/class/module level to identify who owns critical parts of the system (e.g., the database model vs. the frontend UI).
III-B3 Issue Tracker Participation
Issue Tracker Participation captures non-code contributions, which are often overlooked in traditional grading. First, the system tracks who created issues (e.g., bug reports, feature requests) and who commented, resolved, or referenced them in commits. This can be integrated with GitHub Issues, GitLab, Jira, or any other issue tracking system. Additionally, NLP tools parse issue labels to identify the contribution type, such as “documentation,” “performance,” “bug fix,” “enhancement,” and so on. Additionally, the system can collect engagement metrics, such as counting and categorizing interactions (e.g., the number of issues raised/resolved), and measuring responsiveness to team comments or review requests.
III-B4 Code Review Evaluation
Code Review Evaluation is a key component of peer code review, which is critical for software quality and learning. ICA assesses the extent to which each student participates constructively and frequently. The system can conduct the following analysis, if not all, to allow peer code review. The first is Review Frequency, in which the number of pull request reviews and comments per student is tracked, and active participation in feedback loops (suggesting improvements, identifying flaws) is noted. The second one is NLP-based Depth Analysis, which utilizes NLP models (e.g., BERT, RoBERTa) to assess whether a review identifies logical issues, suggests improvements, or references standards or best practices. Here, the surface-level or generic comments receive lower weighting. The third one is Sentiment and Tone Analysis, which detects whether reviews are constructive, neutral, or overly critical/toxic. This encourages professionalism and positive collaboration.
III-B5 Scoring & Visualization
Finally, the scoring and visualization of ICA adopts multimodal aggregation. ICA uses a weighted rubric to combine scores from the four components into a contribution index for each student. Customizable weights enable instructors to emphasize code, communication, or design. A visualization dashboard with contribution heatmaps, pie charts, and timelines helps visualize who contributed to which files, the evolution of each student’s work over time, and participation spikes (e.g., just before deadlines). Furthermore, anomaly detection can be integrated into ICA, enabling AI to flag potential freeloading or over-contributors and to issue team imbalance alerts, prompting early intervention by instructors.
Overall, the Individual Contribution Analyzer (ICA) (as shown in Fig. 2) provides a transparent and data-driven approach to evaluating individual contributions in group projects. By analyzing both technical and collaborative behaviors, ICA promotes fairness, accountability, and reflective learning in collaborative computer science education.
Commit Analysis Code Authorship Issue Tracker and PR Comments Code Review Repository and Communication Artifacts Normalized Individual Contribution Score
Figure 2: Architecture of the Individual Contribution Analyzer (ICA) illustrating how repository and communication artifacts are processed through analytical components to produce the normalized individual contribution score.
III-C Grading Engine (GE)
The Grading Engine (GE) is the final layer of the AI-driven assessment pipeline. It performs weighted aggregation, normalization, anomaly detection, and score interpretation to generate final individual grades from collective group output.
The GE calculates each student’s final score by combining the Project Quality Score (PQS) output from PQAM, which represents the overall technical quality of the group’s work, and the Normalized Contribution Score (NCS) output from ICA, which represents each student’s individual input relative to their peers. The default formula is as follows:
$S_{f}=(PQS\cdot 0.6)+(NCS\cdot 0.4)$ (1)
This formula is configurable. Instructors can adjust the weightings to emphasize team output (e.g., 80/20) or individual effort (e.g., 50/50), depending on pedagogical goals. Grades are automatically generated but flagged for manual review if anomalies (e.g., outliers in contribution) are detected.
To ensure fair grading across teams of varying sizes and dynamics, ICA outputs could be normalized. For example, we can use Z-score Normalization or Min-Max Scaling, in which each student’s raw contribution score is scaled within their team to enable comparison and mitigate the impact of team composition. Also, extremely low contributions (e.g., <10%) may trigger floor limits, and extremely high contributions may be capped to prevent “grade hogging” and support equity. Instructors may optionally award bonuses to exemplary contributors or impose penalties for under-participation based on team evaluations or reflections.
The GE automatically detects irregularities or unfair patterns and flags them for human intervention. For example, it utilizes statistical models (e.g., interquartile range, standard deviation) to identify contributors with significantly higher/lower input than their peers, or teams with extreme PQAM scores but low ICA engagement. Inconsistencies can also be identified, such as mismatches between contribution levels and code authorship (e.g., frequent commits with minimal actual authorship), or imbalanced participation in code reviews or issue resolution. Flagged submissions are queued in an instructor dashboard. The instructor is presented with visualizations of contribution breakdowns, suggested rubric-based adjustments, and space for human notes or overrides.
A key goal of GE is to promote clarity and fairness in the grading process. It will provide breakdown reports for students that show the PQAM score, NCS, and final grade, along with possible visual aids for instructors, such as radar charts and contribution timelines. It also provides feedback prompts that can auto-generate personalized feedback summaries, for example, “Your code coverage and documentation were strong, but code review participation was low,” or “Project quality was excellent, but individual contribution was significantly below team average.” All grading decisions and overrides are documented for transparency in the event of appeals or disputes.
The Grading Engine (GE) acts as the bridge between AI-driven analysis and practical grading outcomes. By combining group quality and individual contribution metrics with anomaly-detection and transparency features, GE ensures that every student receives a fair, evidence-based grade while still allowing instructors to exercise judgment when necessary.
IV System Implementation and Analysis
To validate the effectiveness, fairness, and scalability of the AI-based grading system, a pilot deployment was conducted during a software engineering course at a university in fall 2024. The course consisted of 20 students, divided into five teams, who worked on collaborative projects over 8 weeks. Projects adhered to Agile methodologies and required the use of GitHub for version control, issue tracking, and code submissions. This system was implemented in Python, using Flask for RESTful API development and PostgreSQL for relational data persistence. The frontend dashboards were created using React.js, allowing instructors and students to interact with project metrics, grades, and feedback visually. The key libraries and tools include (1) Scikit-learn, TensorFlow–for building machine learning models (e.g., contribution prediction, NLP classification), (2) Radon, Pylint–static code analysis for metrics such as cyclomatic complexity and style compliance, (3) GitPython–used to interface with Git repositories for extracting commit history and file authorship data, and (4) SpaCy, NLTK–employed in the NLP pipeline to process README files, code comments, and issue tracker text for evaluating documentation and collaboration.
IV-A Workflow Overview
The system followed a modular 5-step pipeline.
IV-A1 Step 1: Project Submission
Students pushed final codebases and documentation to GitHub Classroom, ensuring version control and a consistent submission structure.
IV-A2 Step 2: Data Extraction
Repositories were programmatically mined using the GitHub REST API and GraphQL endpoints to extract commit history and diffs, issue tracker logs, pull requests, code reviews, README files, wikis, and inline comments.
IV-A3 Step 3: Model Evaluation
PQAM evaluated group-level quality using static analysis, test coverage metrics, documentation assessment, and usability checks. ICA parsed Git data and communication records to infer individual contributions using attribution modeling, NLP, and statistical analysis.
IV-A4 Step 4: Grade Calculation
The Grading Engine (GE) combined PQAM and ICA outputs via a configurable formula. Anomaly detection algorithms flagged potential freeloaders or disproportionate contributions for manual review.
IV-A5 Step 5: Dashboard Review
Instructors accessed a secure web portal to view performance breakdowns and individual contributions, adjust or approve final grades, and export results to the LMS.
A system flowchart (Fig. 3) was generated to illustrate the modular interactions from submission through evaluation, grading, and review. The diagram helps stakeholders understand the end-to-end automation and where human judgment is involved.
Step 1: Project Submission Students submit code and documentation via GitHub Classroom Step 2: Data Extraction REST/GraphQL mining – commits, PRs, issues, README Step 3: Model Evaluation PQAM (quality, tests, docs) + ICA (commits, reviews) Step 4: Grade Calculation Combine PQAM + ICA outputs; detect anomalies Anomaly? Manual Review and Adjustment Step 5: Dashboard Review Final review and LMS export YesNo
Figure 3: Workflow of the AI-assisted grading system. Steps 1–4 (blue) form the automated pipeline; anomalies (amber) trigger manual review and dashboard oversight (gray-blue), while normal cases proceed directly.
IV-B Metrics and Evaluation Methods
To assess system validity and impact, three key methods were employed.
IV-B1 Instructor Alignment
AI-generated grades were compared with traditional instructor-assigned grades using the Pearson correlation coefficient (r) to assess consistency.
IV-B2 Student Perception
A post-project survey (N = 53) assessed students’ views on fairness, transparency, and satisfaction using a 5-point Likert scale.
IV-B3 Instructor Effort
Time spent on grading was logged and compared with previous semesters using manual rubrics.
IV-C Pilot Results and Analysis
The AI system demonstrated a strong alignment with human grading decisions, with a Pearson’s r of 0.91, indicating high validity. The Student Satisfaction Survey results show fairness at 4.3/5 and transparency at 4.5/5. Additionally, it demonstrated a 45% reduction in grading effort compared to manual evaluation methods. Finally, 12% of cases required manual intervention (e.g., for flagged anomalies and non-code contributions).
Three anonymized cases illustrate the system’s effectiveness in capturing real-world dynamics. First, contributions were evenly distributed (within ±5%), indicating that AI-assigned grades were uniform and required no manual overrides. Second, one team member authored 58% of the production code and handled most issue resolution. The system elevated their individual score accordingly, with instructor validation. Third, one student inflated activity with frequent whitespace and comment-only commits. The ICA module, using code-diff analysis, downweighted these and flagged the case, thereby avoiding unfair grade inflation.
This pilot result highlights three key advantages of utilizing an AI grading system in collaborative coding projects. First, it demonstrates scalability. It seamlessly managed a class of 20 students and could scale to 200+ with minimal performance degradation. It also demonstrates objectivity by replacing subjective perceptions with data-driven, consistent evaluations. Finally, it proves transparency, as all students (based on an after-class survey) appreciated visibility into how grades were computed, thereby improving trust in the assessment. At the same time, it enhances the incentive to engage because awareness of tracked metrics motivates students to participate consistently.
However, we face particular challenges when we rely entirely on the system to grade. Essential activities, such as design ideation, presentation preparation, and team facilitation, were more strenuous to quantify objectively. Additionally, students who did not use GitHub or issue trackers consistently made fewer contributions. Moreover, students with weaker English skills or non-native grammar patterns were at higher risk of lower documentation scores. Ongoing calibration is needed to avoid bias.
V Ethical Considerations
Integrating AI-assisted grading systems into education requires careful attention to transparency, fairness, privacy, and human oversight. Ethical design ensures that automation enhances rather than undermines trust and equity in assessment. Transparency is foundational to ethical deployment, as students and instructors should understand how the system collects data, evaluates contributions, and calculates final grades. Providing interpretable dashboards, clear rubrics, and documentation of decision logic helps prevent perceptions of “black-box grading” and supports accountability.
Fairness must be actively monitored to avoid unintended bias. NLP-based documentation analysis or behavioral metrics may disadvantage non-native speakers or students with atypical collaboration patterns. Regular audits using fairness metrics, anonymized data during manual reviews, and calibration across cohorts can help maintain equitable grading outcomes. Privacy and data governance are equally critical since repository histories, communication logs, and contribution records contain personally identifiable information. These data must be handled in compliance with institutional and legal standards, such as FERPA and GDPR. They must be supported by informed consent, restricted data access, and well-defined retention and deletion policies.
Finally, human oversight remains essential. The system is designed to augment, not replace, instructor judgment. Manual review tools and structured appeal processes ensure that students retain agency and that instructors can interpret algorithmic outcomes within their pedagogical context. Ethical integration depends on shared responsibility, in which technology provides analytical consistency, and educators ensure empathy and contextual understanding. By embedding transparency, fairness, privacy, and oversight into its design, the proposed system promotes responsible AI use in education. It reinforces the principles of trust and equity that underpin effective learning environments.
VI Future Directions
To extend functionality and improve adaptability, several research and engineering enhancements are planned. Each enhancement is an individual research topic or implementation effort. There is much work ahead; we highlight several primary directions that require further attention.
VI-A IDE Integration
Develop plugins for popular Integrated Development Environments (IDEs), such as Visual Studio Code, IntelliJ, and Eclipse, to track local coding behavior, file ownership, and collaboration patterns. This feature would enable seamless data collection from developers’ working environments, offering insight into how and when students contribute to a shared project. By tracking keystrokes, file changes, commit frequencies, and time spent per file, these plugins can more accurately attribute contributions and detect patterns such as code ownership, multitasking, and collaboration bugs. It also enhances the Individual Contribution Analyzer (ICA) by providing rich metadata beyond Git history. Significant challenges include preserving student privacy, managing data volume, and ensuring cross-platform compatibility.
VI-B Multimedia Artifact Analysis
Utilize machine learning models to evaluate non-code artifacts, including screenshots, wireframes, presentations, and videos. Many software engineering and CS capstone projects include deliverables such as UI mockups, slide decks, and demo videos. Traditional grading overlooks these components or evaluates them manually. By applying computer vision, natural language processing, and audio analysis techniques, the system can automatically assess the clarity, completeness, and professionalism of these artifacts. For example, image recognition could assess consistency in UI wireframes; speech analysis could evaluate clarity and fluency in demo videos. This expansion acknowledges the holistic nature of project work, emphasizing the importance of design and communication skills in addition to coding.
VI-C Enhanced Peer Review
Combine subjective peer feedback with objective ICA metrics to produce more balanced and trustworthy individual assessments. Peer reviews offer a unique perspective into team dynamics and can capture nuances that automated systems miss, such as leadership, initiative, or teamwork quality. However, they can be biased or inconsistent. By integrating peer evaluations with ICA’s behavioral metrics (e.g., code contributions, task tracking, communication patterns), the platform can cross-validate data, flag anomalies, and calibrate scores. Machine learning models may even weight peer input differently based on reviewers’ historical reliability, thereby fostering a more trustworthy and transparent system.
VI-D Adaptive Grading Models
Apply reinforcement learning or similar techniques to dynamically adjust grading models based on past performance and instructor feedback. Current grading systems rely on fixed weights (e.g., 60% project quality, 40% individual contribution). Adaptive grading would instead learn from instructor corrections—such as grade adjustments or override patterns—and iteratively improve the scoring logic. Over time, the system can learn how to handle edge cases, identify scoring inconsistencies, and personalize grading for different course contexts. This reduces manual grading effort and improves grading accuracy, especially in large or diverse classes.
VI-E Cross-Disciplinary Expansion
Adapt the platform for use in disciplines beyond computer science, such as business, design, humanities, and engineering. Many academic fields assign collaborative group projects but lack structured tools for effective assessment. By abstracting core features (e.g., contribution tracking, artifact analysis, peer review), the platform can support project-based learning in various domains. For example, in design, it may evaluate the evolution of creative assets; in business, it could track contributions to a business plan or pitch deck. This would require customizing assessment models, interfaces, and rubrics to align with discipline-specific expectations and workflows.
VI-F Bias Audits
Implement regular fairness assessments to detect and mitigate grading bias across gender, language, or demographic groups. AI-powered assessment systems risk encoding or amplifying biases—especially when evaluating subjective artifacts or using peer input. Conducting systematic audits using fairness metrics (e.g., disparate impact, equal opportunity) helps detect whether certain groups are under- or over-represented among top scorers. This could involve anonymizing submissions during evaluation, monitoring statistical parity, and incorporating fairness-aware ML models. Transparent reporting and instructor tools to review flagged cases would promote ethical and equitable grading.
VI-G LMS & Rubric Integration
Enable the direct export of grades, comments, and rubrics into Learning Management Systems (LMS) such as Canvas, Moodle, Brightspace, and Blackboard. For broad classroom adoption, the system must integrate with existing instructor workflows. By aligning with LMS APIs, the platform can automate the delivery of final grades, individualized feedback, and rubric-based assessments, thereby saving time and reducing data-entry errors. Furthermore, bidirectional integration may allow the system to pull course context, deadlines, and participation data to inform grading. This streamlining is essential for scaling the use of this approach across institutions and ensuring consistency with institutional grading policies.
VII Conclusion
This paper presents a robust AI-powered grading system that addresses key challenges in assessing collaborative group projects in computer science education. By leveraging data from code repositories, communication logs, and automated analysis tools, the system enhances fairness, scalability, and pedagogical insight. Initial trials confirm its efficacy, demonstrating strong alignment with instructor evaluations, reduced grading workload, and improved student trust in assessment transparency. Future expansions will focus on handling non-code contributions, multimedia artifacts, and adaptive grading models applicable across multiple disciplines. These enhancements aim to capture a broader spectrum of collaborative work, ensuring that creativity, communication, and teamwork receive fair recognition alongside technical achievement.
Ultimately, the promise of AI-driven grading lies in collaboration between humans and intelligent systems rather than automation alone. The system’s analytical precision complements the instructor’s judgment and empathy, creating assessments that are both data-informed and human-centered. Guided by transparency and fairness, this partnership can transform grading into a reflective learning process, reshaping how educators evaluate and support learning in project-based environments.